OpenStudy (anonymous):

USE the formula to find the value of the remaining variable. P=2L+2W;L=10,W=78 P=_____

OpenStudy (anonymous):

Just substitute in L and W.

OpenStudy (anonymous):

how would i do that..

OpenStudy (anonymous):

i.e. P=2L+2W P=2(10)+2(78)

OpenStudy (anonymous):

2*10+2(78)=1716 is that correct

OpenStudy (anonymous):

the correct answer is 176

OpenStudy (anonymous):

Yep, its 176 :)

OpenStudy (anonymous):

can u help me with another problem

OpenStudy (anonymous):

The sum of a page number on the facing pages of a book is 89. What are the page numbers. The number on the left page is___ The number onthe right page is____

OpenStudy (anonymous):

Well, lets give the first page number the symbol "x". because the second page number has to be one more than "x", it is "x+1". Therefore, the equation must be...?

OpenStudy (anonymous):

x+1=89

OpenStudy (anonymous):

it would be 2x+1=89 because you are finding the sum of the two pages :)

OpenStudy (anonymous):

So now just solve to find x :)

OpenStudy (anonymous):

x=44

OpenStudy (anonymous):

Yep :) so the first page number is 44, and the second is 45 :)

OpenStudy (anonymous):

Solve for g. A=gr